# Modulate and Reconstruct: Learning Hyperspectral Imaging from Misaligned Smartphone Views

### v2

@@ 2 Related Work
Numerous approaches have aimed to capture multispectral or hyperspectral information without expensive hardware. Early work by Helling et al. [15] employed a grayscale camera with a rotating filter wheel, while Valero et al. [16] used an RGB camera with three interchangeable filters. Oh et al. [14] captured scenes with three different cameras, leveraging the variation in their spectral sensitivities. More recently, Sharma et al. [13] demonstrated that consumer mobile devices with both RGB and NIR sensors can achieve extended spectral capture (400–1000 nm). Although these systems reduce costs, they either involve long capture times or assume no misalignment between successive captures.
Hyperspectral reconstruction from RGB
@@
Traditional HSR methods model spectra using sparse coding [17], dictionary learning [18], or manifold embeddings [19], based on the low-dimensional structure of hyperspectral data and the rarity of metamers [20]. These approaches are computationally efficient but often lack the capacity to incorporate global context from the input image, making them less robust to complex natural scenes. Recent methods based on deep learning have achieved significant advances, particularly those developed through the NTIRE spectral reconstruction challenges [21, 22, 23]. Early deep models used CNNs [9, 24, 25, 26, 27, 28], while newer transformer-based approaches like MST++ [8] and MSFN [29] introduced attention along spectral or spatial dimensions. However, nearly all research relies on synthetic RGB inputs rendered from hyperspectal images (HSIs) using color matching functions (CMFs), assuming perfect alignment and access to camera parameters. These assumptions do not hold in practice, limiting model generalizability. We address these limitations training and evaluating on real-world data with acquisition artifacts and misalignment.
Handling misalignment
@@
Misalignment between inputs and ground truth is a well-known challenge in video and reference-based super-resolution tasks. Optical flow (OF) [30, 31], deformable convolutions [32, 33], and attention mechanisms [34] have been proposed to mitigate this. In our context, misalignment also arises due to ground truth HSIs being not spatially aligned with the RGB input. Zhang et al. [35] addressed this by warping ground truth toward the input using OF, enabling pixel-level evaluation. Elezabi et al. [36] proposed contextual losses and pseudo-aligned inputs as training strategies. We adopt the interpretable and evaluation-friendly approach [35] of warping the ground truth to the input using OF, allowing accurate pixel-wise supervision and metric computation.
Hyperspectral datasets
@@
Several datasets support HSR research. Early datasets like CAVE [37] and Harvard [38] provided controlled hyperspectral measurements. Later datasets such as ICVL [17], KAUST [39], and ARAD_1K [23] focused on enabling data-driven HSR methods. These datasets contain either radiance or reflectance data, but most lack real RGB images and instead simulate RGB via CMFs, which fails to capture the characteristics of camera pipelines. Moreover, they often assume perfect alignment, which does not hold in practical settings. Recent datasets like BeyondRGB [5] and MobileSpec [4] address some of these issues by including real RGB captures. BeyondRGB includes color charts and lightsource spectrum estimation, while MobileSpec offers aligned RGB-HSI pairs. However, they still face trade-offs between alignment, diversity, and color reference availability.We build on these efforts by introducing Doomer dataset with real RGB images, misaligned hyperspectral data, and in-scene color references thus enabling realistic reconstruction under natural capture conditions.
@@ 4 Doomer Dataset
Overview
To support training and evaluation for the MI-HSR task introduced in Sec. 3, we collect a new dataset, which we call Doomer [^1] The name Doomer is inspired by the subcultural aesthetic: most scenes were collected under gloomy or overcast weather conditions, in contrast to the brightly lit existing datasets.. Existing hyperspectral datasets are not suited for our setup: none provide spatially misaligned multi-view RGB observations. Furthermore, no existing dataset aligns with our specific hardware configuration (a triple-camera smartphone with custom spectral filters), making new data collection a necessity.
Doomer contains 143 real-world scenes captured using a Huawei Mate 40 Pro smartphone equipped with Main, Tele, and Wide cameras, along with a Specim IQ hyperspectral camera for ground truth. In each scene, we record three smartphone RAW images, two with custom spectral filters and one unfiltered as well as a 111-band (400 – 730 nm) hyperspectral image. Example captures are shown in Fig. 3.
@@
Capture setup
The data acquisition rig consists of a Specim IQ hyperspectral camera, a Huawei Mate 40 Pro smartphone mounted in a 3D-printed case with slots for spectral filters and a gray reference sphere (VFX ball).
The entire system is mounted on a tripod, with the phone positioned to rotate along a vertical axis for alternating captures (Fig. 4). This design allows the smartphone and hyperspectral camera to image scenes from nearly identical viewpoints, minimizing parallax and occlusion. The gray ball is connected to a rigid rod that allows to regulate the position of the ball in the scene.
@@
Preprocessing pipeline
Each four-image scene group (three RGB + one HSI) undergoes standardized preprocessing (see Supp.). We also normalize field of view and resolution across sensors. Specifically, we estimate pairwise homographies between each RGB image and the Tele view using SIFT keypoints and RANSAC. All images are then warped to the Tele frame and cropped accordingly (Fig. 5). When automatic registration fails, manual alignment is used. Despite this correction, residual geometric misalignment remains due to parallax and non-planar scene structure — motivating our use of alignment-aware HSR models. Finally, all images are downsampled to match the hyperspectral resolution, originally $194\times 259$ cropped to $192\times 256$ for convenience, and spectral grid is resampled to conventional 400–730 nm, $n=34$, to optimize computational costs.
@@ 5.3 Proposed Architecture: MI-MSFN
We propose a revised architecture, Multi-Image MSFN (MI-MSFN), specifically tailored to address input misalignment and heterogeneous spatial content. Here, we reverse the order of modules, applying spatial transformer blocks before spectral ones (Fig. 6). This allows the network to first establish spatial correspondences across the Main, Tele, and Wide inputs thus effectively aligning features by attention before fusing spectral information. We also remove the skip paths to enforce that all feature propagation passes through alignment-aware components.
@@ 6.1 Experimental Setup
Dataset
All the experiments were conducted on our Doomer dataset as it is the only available one for MI-HSR task. Experiments on single-image HSR were only conducted for comparison to MI-HSR, thus featuring the same dataset.
Metrics
We adopt well-known Peak Signal-to-Noise Ratio (PSNR) and Spectral Angle Mapper (SAM). We also introduce Normalized Spectral Error (NSE), which serves as an alternative to MRAE as it does not penalize having few dark bands and reflects the integral nature of radiance:
@@
Implementation details
We split our dataset into train and test subset in proportion 4:1. In all the experiments we used Adam optimizer with learning rate of 0.0004 inherited MSFN (except for AWAN at Sec. 6.3, which showcased the necessity of polynomial scheduler with power of 1.5 starting at learning rate of 0.0001). We trained all the networks on random $64~\times~64$ patches for $10^{4}$ epochs. Every 30 epochs we ran evaluation loop; the final model is the one with the best mean absolute error on the aligned test set. Each training procedure was re-run 10 times with different random seeds, yielding 10 distinct quantitative results. We summarize them by reporting mean and standard deviation.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Code and Data
Click here to download.
A.2 Estimation of Spectral Sensitivity Functions and Transmittance Functions
Spectral sensitivity functions
To estimate SSF of a smartphone camera, we acquired 25 sample pairs of $\mathbf{c}_{i}\in\mathbb{R}^{3}$ and $\mathbf{r}_{i}\in\mathbb{R}^{n}$ corresponding to flat-field illumination (FFI) of narrow-band LED light sources in an integrating sphere. For each LED $i$, we took a photo of FFI (Fig. 8), extracted the central 100×100 patch from it, photometrically normalized it and computed the channel-wise average to get $\mathbf{c}_{i}$. Then we measured the LED radiance spectrum $\mathbf{r}_{i}$ using an X-Rite i1 Pro spectrophotometer with the help of spotread routine from Argyll color management system.
@@ A.2 Estimation of Spectral Sensitivity Functions and Transmittance Functions
Specim IQ calibration
Spectral measurements of X-Rite i1 Pro and Specim IQ are actually inconsistent. It is not suprising: Specim IQ is not designed to be photometrically accurate. Instead, it allows to estimate band-wise relations between the calibration target in a scene and other objects. However, if we estimate the band-wise calibration divisor between X-Rite i1 Pro and Specim IQ, we can futher use it to obtain physically-correct HSIs. We illuminated the integration sphere using a mixture of LEDs that yields even spectral power distribution, obtained X-Rite and Specim IQ measurements of it and divided one by another:
@@ A.3 Spectral Uncertainty
A.4 More on Doomer Dataset
@@
A.5 Rest of the MI-MSFN Architecture
@@
A.6 Additional Experiments on Simulated Data

### v3

@@ 2 Related Work
Low-cost multispectral imaging. Numerous approaches have aimed to capture multispectral or hyperspectral information without expensive hardware. Early work by Helling et al. [15] employed a grayscale camera with a rotating filter wheel, while Valero et al. [16] used an RGB camera with three interchangeable filters. Oh et al. [14] captured scenes with three different cameras, leveraging the variation in their spectral sensitivities. More recently, Sharma et al. [13] demonstrated that consumer mobile devices with both RGB and NIR sensors can achieve extended spectral capture (400–1000 nm). Although these systems reduce costs, they either involve long capture times or assume no misalignment between successive captures.
@@
Hyperspectral reconstruction from RGB. Traditional HSR methods model spectra using sparse coding [17], dictionary learning [18], or manifold embeddings [19], based on the low-dimensional structure of hyperspectral data and the rarity of metamers [20]. These approaches are computationally efficient but often lack the capacity to incorporate global context from the input image, making them less robust to complex natural scenes. Recent methods based on deep learning have achieved significant advances, particularly those developed through the NTIRE spectral reconstruction challenges [21, 22, 23]. Early deep models used CNNs [9, 24, 25, 26, 27, 28], while newer transformer-based approaches like MST++ [8] and MSFN [29] introduced attention along spectral or spatial dimensions. However, nearly all research relies on synthetic RGB inputs rendered from hyperspectal images (HSIs) using color matching functions (CMFs), assuming perfect alignment and access to camera parameters. These assumptions do not hold in practice, limiting model generalizability. We address these limitations training and evaluating on real-world data with acquisition artifacts and misalignment.
@@
Handling misalignment. Misalignment between inputs and ground truth is a well-known challenge in video and reference-based super-resolution tasks. Optical flow (OF) [30, 31], deformable convolutions [32, 33], and attention mechanisms [34] have been proposed to mitigate this. In our context, misalignment also arises due to ground truth HSIs being not spatially aligned with the RGB input. Zhang et al. [35] addressed this by warping ground truth toward the input using OF, enabling pixel-level evaluation. Elezabi et al. [36] proposed contextual losses and pseudo-aligned inputs as training strategies. We adopt the interpretable and evaluation-friendly approach [35] of warping the ground truth to the input using OF, allowing accurate pixel-wise supervision and metric computation.
@@
Hyperspectral datasets. Several datasets support HSR research. Early datasets like CAVE [37] and Harvard [38] provided controlled hyperspectral measurements. Later datasets such as ICVL [17], KAUST [39], and ARAD_1K [23] focused on enabling data-driven HSR methods. These datasets contain either radiance or reflectance data, but most lack real RGB images and instead simulate RGB via CMFs, which fails to capture the characteristics of camera pipelines. Moreover, they often assume perfect alignment, which does not hold in practical settings. Recent datasets like BeyondRGB [5] and MobileSpec [4] address some of these issues by including real RGB captures. BeyondRGB includes color charts and lightsource spectrum estimation, while MobileSpec offers aligned RGB-HSI pairs. However, they still face trade-offs between alignment, diversity, and color reference availability.We build on these efforts by introducing Doomer dataset with real RGB images, misaligned hyperspectral data, and in-scene color references thus enabling realistic reconstruction under natural capture conditions.
@@ 4 Doomer Dataset
Overview. To support training and evaluation for the MI-HSR task introduced in Section 3, we collect a new dataset, which we call Doomer [^1] The name Doomer is inspired by the subcultural aesthetic: most scenes were collected under gloomy or overcast weather conditions, in contrast to the brightly lit existing datasets.. Existing hyperspectral datasets are not suited for our setup: none provide spatially misaligned multi-view RGB observations. Furthermore, no existing dataset aligns with our specific hardware configuration (a triple-camera smartphone with custom spectral filters), making new data collection a necessity.
Doomer contains 143 real-world scenes captured using a Huawei Mate 40 Pro smartphone equipped with Main, Tele, and Wide cameras, along with a Specim IQ hyperspectral camera for ground truth. In each scene, we record three smartphone RAW images, two with custom spectral filters and one unfiltered as well as a 111-band (400 – 730 nm) hyperspectral image. Example captures are shown in Figure 3.
@@
Capture setup. The data acquisition rig consists of a Specim IQ hyperspectral camera, a Huawei Mate 40 Pro smartphone mounted in a 3D-printed case with slots for spectral filters and a gray reference sphere (VFX ball).
The entire system is mounted on a tripod, with the phone positioned to rotate along a vertical axis for alternating captures (Figure 4). This design allows the smartphone and hyperspectral camera to image scenes from nearly identical viewpoints, minimizing parallax and occlusion. The gray ball is connected to a rigid rod that allows to regulate the position of the ball in the scene.
@@
Preprocessing pipeline. Each four-image scene group (three RGB + one HSI) undergoes standardized preprocessing (see Supp.). We also normalize field of view and resolution across sensors. Specifically, we estimate pairwise homographies between each RGB image and the Tele view using SIFT keypoints and RANSAC. All images are then warped to the Tele frame and cropped accordingly (Fig. 5). When automatic registration fails, manual alignment is used. Despite this correction, residual geometric misalignment remains due to parallax and non-planar scene structure — motivating our use of alignment-aware HSR models. Finally, all images are downsampled to match the hyperspectral resolution, originally $194\times 259$ cropped to $192\times 256$ for convenience, and spectral grid is resampled to conventional 400–730 nm, $n=34$, to optimize computational costs.
@@ 5.3 Proposed Architecture: MI-MSFN
We propose a revised architecture, Multi-Image MSFN (MI-MSFN), specifically tailored to address input misalignment and heterogeneous spatial content. Here, we reverse the order of modules, applying spatial transformer blocks before spectral ones (Figure 6). This allows the network to first establish spatial correspondences across the Main, Tele, and Wide inputs thus effectively aligning features by attention before fusing spectral information. We also remove the skip paths to enforce that all feature propagation passes through alignment-aware components.
@@ 6.1 Experimental Setup
Dataset. All the experiments were conducted on our Doomer dataset as it is the only available one for MI-HSR task. Experiments on single-image HSR were only conducted for comparison to MI-HSR, thus featuring the same dataset.
Metrics. We adopt well-known Peak Signal-to-Noise Ratio (PSNR) and Spectral Angle Mapper (SAM). We also introduce Normalized Spectral Error (NSE), which serves as an alternative to MRAE as it does not penalize having few dark bands and reflects the integral nature of radiance:
@@
Implementation details. We split our dataset into train and test subset in proportion 4:1. In all the experiments we used Adam optimizer with learning rate of 0.0004 inherited MSFN (except for AWAN at Sec. 6.3, which showcased the necessity of polynomial scheduler with power of 1.5 starting at learning rate of 0.0001). We trained all the networks on random $64~\times~64$ patches for $10^{4}$ epochs. Every 30 epochs we ran evaluation loop; the final model is the one with the best mean absolute error on the aligned test set. Each training procedure was re-run 10 times with different random seeds, yielding 10 distinct quantitative results. We summarize them by reporting mean and standard deviation.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Estimation of Spectral Sensitivity Functions and Transmittance Functions
Spectral sensitivity functions. To estimate SSF of a smartphone camera, we acquired 25 sample pairs of $\mathbf{c}_{i}\in\mathbb{R}^{3}$ and $\mathbf{r}_{i}\in\mathbb{R}^{n}$ corresponding to flat-field illumination (FFI) of narrow-band LED light sources in an integrating sphere. For each LED $i$, we took a photo of FFI (Fig. 8), extracted the central 100×100 patch from it, photometrically normalized it and computed the channel-wise average to get $\mathbf{c}_{i}$. Then we measured the LED radiance spectrum $\mathbf{r}_{i}$ using an X-Rite i1 Pro spectrophotometer with the help of spotread routine from Argyll color management system.
@@ A.1 Estimation of Spectral Sensitivity Functions and Transmittance Functions
Specim IQ calibration. Spectral measurements of X-Rite i1 Pro and Specim IQ are actually inconsistent. It is not suprising: Specim IQ is not designed to be photometrically accurate. Instead, it allows to estimate band-wise relations between the calibration target in a scene and other objects. However, if we estimate the band-wise calibration divisor between X-Rite i1 Pro and Specim IQ, we can futher use it to obtain physically-correct HSIs. We illuminated the integration sphere using a mixture of LEDs that yields even spectral power distribution, obtained X-Rite and Specim IQ measurements of it and divided one by another:
@@ A.2 Spectral Uncertainty
A.3 More on Doomer Dataset
@@
A.4 Rest of the MI-MSFN Architecture
@@
A.5 Additional Experiments on Simulated Data